I need the problem factored. Or simply solved for x. 4x^2+12x+36=0
Answer by stanbon(75887)

4x^2+12x+36=0
4(x^2+3x+9) = 0
x^2 + 3x + 9 = 0
x = [-3 +- sqrt(3^2 -4*1*9)]/2
x = [-3 +- sqrt(9-36)]/2
x = [-3 +- sqrt(-27)]/2
x = [-3 +- 3isqrt(3)]/2
===========================
Cheers,
Stan H.